OpenStudy (anonymous):

Find the absolute extrema as well as values of x wherre they occur using the given function and domain f(x)=x+e^(-5x), [-2,2] PLEASE HELP MEEE

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@satellite73

OpenStudy (gorv):

what u do to find xtrema???

OpenStudy (gorv):

u know how to find xtrema???

OpenStudy (gorv):

@aaaaagirl

OpenStudy (anonymous):

\[f'(x)=1-5e ^{-5x},f'(x)=0 gives e ^{-5x}=\frac{ 1 }{ 5 }\] \[e ^{5x}=5,x=\frac{ 1 }{5 }\ln 5\] \[find f(x) at x=-2,2,\frac{ 1 }{ 5 }\ln 5\] see where it is maximum .